# MOSS: Motion-based 3D Clothed Human Synthesis from Monocular Video

### v1

@@ Abstract
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.
@@ Introduction
Recently, 3D Gaussian Splatting(3DGS)  [9, 3, 23, 38, 22, 4, 27] has greatly improved the efficiency and quality of 3D reconstruction by sampling object geometry through Gaussian distributions. Existing work on reconstructing the human body with 3DGS [15, 25] introduces the SMPL as the human body prior, which can restore a more realistic human body. However, these works neglect the hierarchical structure constraints of the kinematic tree [11] and the constraints of global motion information in reconstructed human surface, resulting in blurred joint details. Besides, the restored surface deformation is still underexplored. Another line of works [48, 49] developed normals to model clothing to characterize the deformation superficially. The influence of mutual occlusion between clothing folds on the fold rendering was not considered, resulting in the loss of clothing folds on the reconstructed human surface.
@@
First, with the aim of deformation reconstruction, we propose KGAS, a module that extracts the human body from decomposing the matrix-Fisher distribution parameter, obtaining the principal axis concentration and rotation factor to explicitly control Gaussians for 3DGS rendering of human surface deformation. During the Gaussian layout process, the principal axis concentration performs as a density factor that corrects the sampling probability of Gaussian split to obtain surface deformation-aware Gaussians. In the following split control, the principal axis concentration and the rotation factor dynamically adjust the direction and scale of the Gaussian, realizing that each Gaussian expresses the clothed human body under the constraint of global motion information.
In single-view videos, local deformations often obscure each other, resulting in insufficient supervision of pixel features in 2D images. Consequently, the local deformations may fail to be reconstructed. To address this issue, it is imperative to consider mitigating the problem of local occlusion by capturing local geometrical deformations. We introduce the UID, where the core lies in the covariance matrix, which characterizes the heterogeneity of the spatial structure of the local point cloud. This characterization facilitates the discernment of local density and morphological differences within the point cloud. UID is grounded in explicit human structure retrieval from the surface, enabling localization to areas exhibiting drastic deformations. Subsequently, geometric reconstruction is performed to compensate for these deformations, thereby facilitating the restoration of occluded local geometry.
@@ 1 Related Work
1.2 Human Reconstruction.
In the field of human modeling, many methods have been used to model human shape and pose. One of the commonly used methods is SMPL [28], which is a vertex-based additive human body model that uses linear algebraic models to model shape and pose separately. At this stage, based on 3DGS, some methods [2, 5, 7, 29] use an explicit human structure SMPL priori to aid in the 3DGS optimization process to reconstruct diverse and realistic 3D human models with high-quality. PSAvatar [52] uses discrete geometric primitives to create parametric deformable shape models that are represented in fine detail and rendered with high fidelity with the help of 3D Gaussian. Human101 [25] uses human-centered forward in the SMPL prior to deforming 3D human models. Gaussian animation method to deform the parameters of 3D Gaussian and improve the rendering speed. GauHuman [15] not only introduces SMPL prior, but also further accelerates the 3DGS by introducing KL dispersion. These methods can restore a more realistic human appearance. However, they lack the constraints of the SMPL kinematic tree when considering human joint linkage effects and surface deformation, which can lead to problems such as blurred joint edges or confusion between joints and clothing when reconstructing complex human poses or when there is clothing occlusion.
1.3 Surface Reconstruction.
@@ 2 Method
Overview. In this paper, we aim to efficiently reconstruct high-quality clothed humans from single-view videos involving big movements. Our approach builds on the assumption that the corresponding SMPL parameters $\beta$, $\theta$, the calibrated camera parameters, and the human region mask are all given. Figure 2 illustrates our framework. We first use joint rotations to control the spatial position and orientation during the 3D Gaussian density transformation (Section 2.2, Section 2.3). The density of the Gaussian distribution is captured by the matrix-Fisher distribution, and the transformation of Gaussians is guided by density. Additionally, significant surface deformation is identified by analyzing the direction of maximum surface geometric variations (Section 2.4).). MOSS distinguishes between the human body and its surface, and constrains the newly generated human body surface Gaussian with whole-body motion information. This enables the reconstruction of more realistic and reasonable details of human surface deformation.
@@ 2.1 Preliminary
SMPL and LBS Transformation. The SMPL human model comprises a human skeleton with $N=6890$ vertices and $k=24$ joints. In this context, let’s define the joint rotation as $\theta_{joint}=[\theta_{joint_{0}}^{T},...,\theta_{joint_{K}}^{T}]^{T}$, where $K=23$, and $\theta_{joint_{k}}\in R^{3}$ denotes the rotation of part $k$ relative to its parent in the kinematic tree. Additionally, $\theta_{joint_{0}}$ refers to the orientation of the root part.
With Linear Blend Skinning(LBS), the position $p^{\prime}_{i}$ of the $i$th vertex $p_{i}$ in canonical space after skinning is denoted as: ${p}_{i}^{\prime}=\sum\nolimits_{k=1}^{K}w_{ki}G_{0k}(\theta_{joint},J(\beta)){p}_{i}+G_{1k}(\theta_{joint},J(\beta))$, where $G_{0k}(\theta_{joint},J(\beta))$ is the rotation transformation matrix of the joint $k$ that depends on the joint rotation $\theta_{joint}$ and the pose change parameter $\beta$ of human. Similarly, $G_{1k}(\theta_{joint},J(\beta))$ is the translation transformation matrix. $w_{ki}$ is the skinning weight of the joint $k$ on vertex $i$, which determines the extent to which joint $k$ affects vertex $i$.
Gaussian Splatting. We use $f_{\mathrm{3DGS}}=\sum\nolimits_{i}G(x,y,z,\mu_{i},\Sigma_{i})\cdot c_{i}$ to express 3DGS, where $c_{i}$ denotes color. where $G$ is a Gaussian function with inputs as a center (position) $x,y,z$, mean $\mu_{i}$ and covariance $\sum\nolimits_{i}$:
$G(x)=e^{-\frac{1}{2}x^{T}\Sigma^{-1}x},\Sigma=R_{g}S_{g}^{T}S_{g}R_{g}^{T}.$ (1)
Matrix-Fisher distribution. All 3D rotation matrices lie on a nonlinear and closed manifold SO(3). The matrix-Fisher distribution is defined over SO(3) [45, 20, 8]. For a given random rotation matrix $\textbf{R}\in SO(3)$, 3D special orthogonal groups can be defined as $SO(3)=\{\textbf{R}\in\mathbb{R}^{3\times 3}|=\textbf{R}^{T}\textbf{R}=\textbf{I},det(\textbf{R})=1\}$. The matrix-Fisher distribution is obtained according to the matrix rotation parameter denoted as $\theta_{joint_{i}}$. The probability density function of $R$ is given by the matrix-Fisher distribution:
$p(\textbf{R}|\theta_{joint_{i}})=\frac{1}{c(\theta_{joint_{i}})}\exp\left(\text{tr}\left(\theta_{joint_{i}}^{T}\textbf{R}\right)\right)=\mathcal{M}(\textbf{R};\theta_{joint_{i}}).$ (2)
@@
2.2 Global Motion Extraction
To restore the joint details of the clothed human body, one of the challenges is to incorporate the SMPL kinematic tree. The SMPL kinematic tree constrains the joint changes of the human body. To this end, we guide the 3D Gaussians with global motion, the source of surface deformation, from SMPL kinematic tree, which constraints the new Gaussians precisely and to the finest granularity (Section 2.3). In this section, our first use of principal axis concentration matrices linked to scale transformations of Gaussian distributions is a breakthrough in the exploration of the semantics of Fisher distributions.
Matrix-Fisher over the SMPL kinematic tree. Implementing 3DGS for clothed humans, the Gaussians are typically weakly constrained by the entire body. We first aim to extract joint motion information. Inspired by  [24, 33, 41], we attempt to utilize the joint rotation distribution to extract global motion and subsequently employ it to guide the 3D Gaussian density transformation.
We perform a proper singular value decomposition $\mathbf{\theta_{joint}}$ of the matrix parameters $\mathbf{\theta_{joint}}$ of the joint rotation distribution $\mathbf{\theta_{joint}}=\boldsymbol{U}\boldsymbol{S}\boldsymbol{V}^{T}$. where $\boldsymbol{U}$ and $\boldsymbol{V}$ are orthogonal matrices, $U$ characterizes the principal axes, and $S=diag(s_{1},s_{2},s_{3})$ is a diagonal matrix consisting of the singular values $s_{1},s_{2},s_{3}$ of $\mathbf{\theta_{joint}}$ $s_{1}>s_{2}>s_{3}$ as a diagonal matrix. The singular values can dominate the degree of concentration of the distribution along the principal axes. It is considered a scaling factor that affects the scaling matrix $S_{g}$ of the Gaussian distribution.
The most probable rotation $R_{mode}$ of the joints is obtained by matrix-Fisher’s mode (Section 2.1), which contains joint motion information. To make new Gaussians correspond with the SMPL kinematic tree’s comovement, we use $R_{mode}$ to provide direction guidance for the rotation $R_{g}$ of the new 3D Gaussians.
Thus, we complete the human joint motion information extraction, obtaining the two factors $S$ and $R_{mode}$ to be used in Section 2.3.
Joint-Driven Orientation Refinement. After defining the matrix-Fisher over the SMPL kinematic tree, we further extract the global motion information during human pose transformation and utilize it to guide the displacement of Gaussians. We propose Joint-Driven Orientation Refinement(JNTM), which learns the correlation weights of joints on the influence of all Gaussians, and explicitly controls the Gaussians to realize ”pulling one to move the whole body”. We use Joint Driver in JNTM to propagate motion information to every Gaussian. Technically, we implement Joint Driver via cross-attention for the sake of simplicity and directness.
To characterize the joint linkage information of the whole body, we introduce the grandfather node relationship of the kinematic tree to constrain the matrix-Fisher distribution of each joint. We model the matrix-Fisher distribution of the rotation matrices of the child nodes of a kinematic tree after the rotation matrices of its parent nodes are known: $\mathcal{M}\left(\theta_{joint_{i}};R_{i}\right)=p(R_{i}|\theta_{joint_{j}},j\in p(i)),$ where $\theta_{joint_{i}}$ denotes the $i$th joint rotation parameter to be estimated. $\theta_{joint_{i}}$ can be estimated more exactly by combining the kinematic tree and introducing the prior information of the parent nodes in $p(i)$ by autoregression. The estimated child node $\theta_{joint_{i}}=f_{i}(\theta_{joint_{i}},\{\theta_{joint_{j}}\}_{j\in P(i)})$.
The matrix-Fisher over the SMPL kinematic tree dynamically represents the real-time global motion which serves as guidance for new Gaussians. The parameters of the matrix-Fisher distribution are estimated by autoregression.
2.3 Motion Guidance in 3D Gaussian Density Control
@@
When performing large-scale motion, the relative displacement of limbs is usually larger than that of the trunk. However, Gaussians learning the limb-aroused cloth folds around the trunk are produced by human trunk Gaussians, lacking essential global motion information. Hence, we leverage global motion extracted from Section 2.2 to guide the density transformation process of Gaussians for better capture of the human body under substantial motion.
@@
Density Perceptual Sampler. The scaling factor of 3DGS is corrected with the principal axis concentration matrix $S$ to change the size of the sampling area in order to make it closer to the real surface shape. This contributes to a more accurate representation of the surface deformation by the Gaussians. The corresponding Gaussian distribution obtained after the adjustment by $S$ is $g(x)=e^{-\frac{1}{2}x^{T}\Sigma^{\prime-1}x}$, as shown in Figure 3. The displacement of the new Gaussian denoted as $\Delta\textbf{x}$ is generated from $g(x)$, and the position of the center of the new Gaussian ellipsoid $\textbf{x}_{new}$ is obtained from the original Gaussian by this displacement transformation: $\textbf{x}_{new}=\Delta\textbf{x}+\textbf{x}_{source},$ so the new Gaussian obtained by cloning from the original is $G_{new}=G(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\Sigma^{-1}x_{new}}.$ The locations of the resulting new Gaussians are closer to the surface, allowing these Gaussians to more realistically represent the geometric deformation of the surface.
@@
Motion Guidance for Gaussians. Next, we use the $S$ and $R_{mode}$, which contain global motion information, to control the density transformation process of the Gaussians. The rotation probability concentration matrix $S$ characterizing the different principal axes is used as a scale correction factor for the Gaussian to control the size of the Gaussians. The joint rotation state estimation $R_{mode}$ is used to control the direction of rotation of the Gaussians. matrix-Fisher extracts real-time motion information that is used to guide the Gaussians to learn human posture changes in real-time. In the Gaussian clone and split process, the estimated $S$ and $R_{mode}$ guide the scaling and rotation of the new Gaussians obtained after cloning, respectively: $S_{g}^{\prime}=mat\_multi(S,S_{g})$, $R_{g}^{\prime}=mat\_multi(R_{mode},R_{g}).$
@@
$g(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\Sigma^{\prime-1}x_{new}}$ and $\Sigma^{\prime}=R_{g}S_{g}^{\prime T}S^{\prime}_{g}R_{g}^{T}$ are obtained, corresponding to the Gaussian distribution. Consequently, KGAS (Section 2.2, Section 2.3) achieves Gaussians density perception. Besides, the new Gaussians can be effectively constrained by the whole-body motion, leading to the generation of more realistic and plausible human body and surface details.
2.4 Surface Deformation Detector

### v2

@@ Abstract
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding Author.This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) [9, 3, 23, 41, 22, 4, 28] has greatly improved the efficiency and quality of 3D reconstruction by sampling object geometry through Gaussian distributions. Existing work on reconstructing the human body with 3DGS [15, 26] introduces the SMPL [29] as the human body prior, which can restore a more realistic human body. However, these works neglect the hierarchical structure constraints of the kinematic tree [11] and the constraints of global motion information in reconstructed human surface, resulting in details missing in the joints of the moving human body and the surface parts of tight-fitting clothing. However, the restored surface deformation is still underexplored. Another line of works [52, 53] developed normals to model clothing to characterize the deformation superficially. The influence of mutual occlusion between clothing folds on the fold rendering was not considered, resulting in the loss of clothing folds on the reconstructed human surface.
@@
To effectively tackle the issue of detail blurring from human body movement, we propose KGAS. This module links global motion with Gaussian point shapes, using the matrix-Fisher [48, 20, 8] probability distribution of human body joints to derive rotation factors and principal axis concentrations. The body pose is parameterized by relative 3D joint rotations along the SMPL kinematic tree, which can be represented using rotation matrices. Regressing rotation matrices using neural networks is challenging because they belong to SO(3) [24, 36], a non-linear 3D manifold with a different topology, where unconstrained neural network outputs are located. However, it is possible to define probability density functions over the Lie group SO(3), such as the matrix-Fisher distribution, whose parameter is an element which can be easily regressed with a neural network [35, 44]. We use a hierarchical probability distribution over relative 3D joint rotations along the SMPL kinematic tree, where the probability density function of each joint’s relative rotation matrix is a matrix-Fisher distribution conditioned on the parents of that joint in the kinematic tree. We train a deep neural network to predict the parameters of such a distribution over body pose, along with a Gaussian distribution over SMPL shape. By globally associating these factors and controlling Gaussians for 3DGS rendering, we achieve precise surface depiction. The principal axis concentration in the Gaussian layout process adjusts sampling probabilities for capturing surface deformations accurately. Through dynamic adjustments in direction and scale, the rotation factor and principal axis concentration ensure each Gaussian faithfully represents the human body’s surface deformation under global motion constraints.
Expanding on KGAS, we introduce UID to tackle clothing wrinkle loss caused by limited 2D image supervision due to local occlusion. By evaluating the degree of directional change in the local distribution of Gaussians through comparing the angles between the normal vectors of adjacent Gaussian points with a set threshold, UID can identify and densely process regions on the surface that exhibit significant deformation. In these regions, we employ the KGAS method for geometric reconstruction compensation. This facilitates the restoration of obscured structural details.
@@ 2 Related Work
2.2 Human Reconstruction
In the field of human modeling, many methods have been used to model human shape and pose. One commonly used method is SMPL, which is a vertex-based additive human body model that uses linear algebraic models to separately model shape and pose. The SMPL model employs Linear Blend Skinning (LBS) [25] technology to smoothly deform the 3D mesh based on the influence of multiple bones in a skeletal animation system, thereby achieving efficient and realistic human pose modeling.
At this stage, based on 3DGS, some methods [2, 5, 7, 30] use an explicit human structure SMPL priori to aid in the 3DGS optimization process to reconstruct diverse and realistic 3D human models with high-quality. PSAvatar [56] uses discrete geometric primitives to create parametric deformable shape models that are represented in fine detail and rendered with high fidelity with the help of 3D Gaussian. Human101 [26] uses human-centered forward in the SMPL prior to deforming 3D human models. Gaussian animation method to deform the parameters of 3D Gaussian and improve the rendering speed. GauHuman [15] not only introduces SMPL prior, but also further accelerates the 3DGS by introducing KL dispersion. These methods can restore a more realistic human appearance. However, they lack the constraints of the SMPL kinematic tree when considering human joint linkage effects and surface deformation, which can lead to problems such as blurred joint edges or confusion between joints and clothing when reconstructing complex human poses or when there is clothing occlusion.
2.3 Surface Reconstruction
@@ 3 Method
3.1 Overview
In this paper, we aim to efficiently reconstruct high-quality clothed humans from single-view videos involving big movements. Our approach builds on the assumption that the corresponding SMPL parameters $\beta$, $\theta$, the calibrated camera parameters, and the human region mask are all given. Figure 2 illustrates our framework. We first introduce KGAS which used joint rotations to control the spatial position and orientation during the 3D Gaussian density transformation (Section 3.3). The density of the Gaussian distribution is captured by the matrix-Fisher distribution, and the transformation of Gaussians is guided by density. Then, we propose UID to accurately capture the Gaussian points at corresponding positions based on the direction of the greatest geometric deformation of the reconstructed object’s surface. (Section 3.4). MOSS differentiates between the human body and its surface, and constrains the newly generated surface Gaussian of the human body using whole-body motion information. This approach allows for the reconstruction of more realistic and accurate details of human surface deformation.
@@ 3.2 Preliminary
3.2.1 SMPL and LBS transformation
The SMPL human model comprises a human skeleton with $N=6890$ vertices and $k=24$ joints. In this context, let’s define the joint rotation as $\theta_{joint}=[\theta_{joint_{0}}^{T},...,\theta_{joint_{K}}^{T}]^{T}$, where $K=23$, and $\theta_{joint_{k}}\in R^{3}$ denotes the rotation of part $k$ relative to its parent in the kinematic tree. Additionally, $\theta_{joint_{0}}$ refers to the orientation of the root part.
With LBS, the position $p^{\prime}_{i}$ of the $i$th vertex $p_{i}$ in canonical space after skinning is denoted as:
${p}_{i}^{\prime}=\sum\nolimits_{k=1}^{K}w_{ki}G_{0k}(\theta_{joint},J(\beta)){p}_{i}+G_{1k}(\theta_{joint},J(\beta)),$ (1)
where $G_{0k}(\theta_{joint},J(\beta))$ is the rotation transformation matrix of the joint $k$ that depends on the joint rotation $\theta_{joint}$ and the pose change parameter $\beta$ of human. Similarly, $G_{1k}(\theta_{joint},J(\beta))$ is the translation transformation matrix. $w_{ki}$ is the skinning weight of the joint $k$ on vertex $i$, which determines the extent to which joint $k$ affects vertex $i$.
3.2.2 Matrix-Fisher distribution
All 3D rotation matrices lie on a nonlinear and closed manifold SO(3). The matrix-Fisher distribution is defined over SO(3). For a given random rotation matrix $\textbf{R}\in SO(3)$, 3D special orthogonal groups can be defined as $SO(3)=\{\textbf{R}\in\mathbb{R}^{3\times 3}|=\textbf{R}^{T}\textbf{R}=\textbf{I},det(\textbf{R})=1\}$. The matrix-Fisher distribution is obtained according to the matrix rotation parameter denoted as $\theta_{joint_{i}}$. The probability density function of $R$ is given by the matrix-Fisher distribution:
$\mathcal{M}(\textbf{R};\theta_{joint_{i}})=p(\textbf{R}|\theta_{joint_{i}})=\frac{1}{c(\theta_{joint_{i}})}\exp\left(\text{tr}\left(\theta_{joint_{i}}^{T}\textbf{R}\right)\right).$ (2)
@@ 3.2.2 Matrix-Fisher distribution
3.2.3 3D Gaussian Splatting
We use $f_{\mathrm{3DGS}}=\sum\nolimits_{i}G(x,y,z,\mu_{i},\Sigma_{i})\cdot c_{i}$ to express 3DGS, where $c_{i}$ denotes color. where $G$ is a Gaussian function with inputs as a center (position) $x,y,z$, mean $\mu_{i}$ and covariance $\sum\nolimits_{i}$:
$G(x)=e^{-\frac{1}{2}x^{T}\boldsymbol{\Sigma}^{-1}x},\boldsymbol{\Sigma}=R_{g}S_{g}^{T}S_{g}R_{g}^{T}.$ (4)
3.3 Kinematic Gaussian Locating Splatting(KGAS)
@@
KGAS performs 3DGS based on global motion information. In this subsection, we will explain how to extract global motion information and how it influences 3DGS, respectively.
3.3.1 Global motion extraction
Implementing 3DGS for clothed humans, the Gaussians are typically weakly constrained by the entire body. We attempt to extract global motion and subsequently employ it to guide the 3D Gaussian density transformation.
To obtain global motion information, we first need to extract local motions. Inspired by [24, 36, 44], We perform a proper singular value decomposition $\mathbf{\theta_{joint}}$ of the matrix parameters $\mathbf{\theta_{joint}}$ of the joint rotation distribution $\mathbf{\theta_{joint}}=\boldsymbol{U}\boldsymbol{S}\boldsymbol{V}^{T}$. where $\boldsymbol{U}$ and $\boldsymbol{V}$ are orthogonal matrices, $U$ characterizes the principal axes, and $\boldsymbol{S}=diag(s_{1},s_{2},s_{3})$ represents a diagonal matrix comprising the singular values $s_{1}>s_{2}>s_{3}$ of $\mathbf{\theta_{joint}}$. The singular values dominate the degree of concentration of the distribution along the principal axes, so we use $\boldsymbol{S}$ as evidence to represent the magnitude of directional changes in each joint’s movement.
The most probable rotation $R_{mode}$ of the joints is obtained by matrix-Fisher’s mode (Section 3.2), which contains joint motion information. We utilize $R_{mode}$ to provide directional guidance for the movement of each joint.
Thus, we complete the local human joint motion information extraction. To capture the joint linkage information of the entire body, we introduce the node relationships within the kinematic tree to constrain the matrix-Fisher distribution for each joint. We refer to this approach as Joint-Driven Orientation Refinement (JNTM). We propose to model the matrix-Fisher distribution of the rotation matrices associated with the child nodes of a kinematic tree, given the rotation matrices of their parent nodes. Specifically, we express this as:
$\mathcal{M}\left(\theta_{joint_{i}};R_{i}\right)=p(R_{i}|\theta_{joint_{j}},j\in p(i)),$ (5)
where $\theta_{joint_{i}}$ denotes the $i$th joint rotation parameter to be estimated. $\theta_{joint_{i}}$ can be estimated more exactly by combining the kinematic tree and introducing the prior information of the parent nodes in $p(i)$ by autoregression. The estimated child node $\theta_{joint_{i}}$ is represented as:
$\theta_{joint_{i}}=f_{i}(\theta_{joint_{i}},\{\theta_{joint_{j}}\}_{j\in P(i)}).$ (6)
The matrix-Fisher distribution over the SMPL kinematic tree dynamically represents real-time global motion, providing valuable guidance for Gaussian Splatting. Subsequently, we will illustrate how this global motion guides 3D Gaussian Splatting in the following subsection.
3.3.2 Motion-Based 3D Gaussian Splatting
@@
When executing large-scale motion, the relative displacement of limbs typically exceeds that of the trunk. However, the Gaussian models that capture cloth folds around the trunk, induced by limb movements, are derived solely from the Gaussians representing the human trunk. Consequently, these models lack crucial global motion information. Therefore, we utilize the global motion extracted in the preceding subsection to guide the density transformation process of the Gaussians, aiming to enhance the representation of the human body during significant motion.
@@
To adjust the size of the sampling area and make it more representative of the actual surface shape, the scaling factor of 3DGS is corrected using the principal axis concentration matrix $\boldsymbol{S}$. We refer to this method as the Density Perceptual Sampler. This contributes to a more accurate representation of the surface deformation by the Gaussians. As shown in Figure 3, The corresponding Gaussian distribution $G^{\prime}(x)$ obtained after the adjustment by $\boldsymbol{S}$ is represented as:
$G^{\prime}(x)=e^{-\frac{1}{2}x^{T}\boldsymbol{\Sigma}^{-1}x},$ (7)
where $\boldsymbol{\Sigma}=R_{g}\boldsymbol{S}^{T}\boldsymbol{S}R_{g}^{T}$. The displacement of the new Gaussian denoted as $\Delta\textbf{x}$ is generated from $G^{\prime}(x)$, and the position of the center of the new Gaussian ellipsoid $\textbf{x}_{new}=\Delta\textbf{x}+\textbf{x}_{source}$ is obtained from the original Gaussian by this displacement transformation. So the new Gaussian $G_{clone}$ obtained by cloning from the original is represented as:
$G_{clone}=G(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\boldsymbol{\Sigma}^{-1}x_{new}}.$ (8)
The locations of the resulting new Gaussians are closer to the surface, allowing these Gaussians to more realistically represent the geometric deformation of the surface.
@@
Next, we use the $\boldsymbol{S}$ and $\boldsymbol{R_{mode}}$, which contain global motion information, to control the density transformation process of the Gaussians. The rotation probability concentration matrix $\boldsymbol{S}$ characterizing the different principal axes is used as a scale correction factor for the Gaussian to control the size of the Gaussians. The joint rotation state estimation $\boldsymbol{R_{mode}}$ is used to control the direction of rotation of the Gaussians. matrix-Fisher extracts real-time motion information that is used to guide the Gaussians to learn human posture changes in real-time. In the Gaussian clone and split process, the estimated $\boldsymbol{S}$ and $\boldsymbol{R_{mode}}$ guide the scaling and rotation of the new Gaussians obtained after cloning, respectively:
$\boldsymbol{S_{g}}^{\prime}=mat\_multi(\boldsymbol{S},\boldsymbol{S_{g}}),\boldsymbol{R_{g}}^{\prime}=mat\_multi(\boldsymbol{R_{mode}},\boldsymbol{R_{g}}).$ (9)
The final Gaussian:
$G(x_{new})=e^{-\frac{1}{2}x_{new}^{T}\boldsymbol{\Sigma^{\prime}}^{-1}x_{new}},$ (10)
where $\boldsymbol{\Sigma^{\prime}}=\boldsymbol{R^{\prime}_{g}}\boldsymbol{S^{\prime}}^{T}\boldsymbol{S^{\prime}}\boldsymbol{R^{\prime}_{g}}^{T}$ are obtained, corresponding to the Gaussian distribution.
@@
Consequently, KGAS achieves Gaussian density perception with the guidance of global motion. Furthermore, the integration of whole-body motion effectively constrains the new Gaussians, thereby enhancing the generation of realistic human body and surface details.
3.4 Surface Deformation Detector(UID)